[The following information applies to the questions displayed below.]

Matt and Meg Comer are married and file a joint tax return. They do not have any children. Matt works as a history professor at a local university and earns a salary of $68,000. Meg works part time at the same university. She earns $33,000 a year. The couple does not itemize deductions. Other than salary, the Comers’ only other source of income is from the disposition of various capital assets (mostly stocks). (Use the tax rate schedules, Dividends and Capital Gains Tax Rates.) (Round your final answers to the nearest whole dollar amount.)

a. What is the Comers’ tax liability for 2020 if they report the following capital gains and losses for the year?

Short-term capital gains $ 9,000
Short-term capital losses (2,000 )
Long-term capital gains 15,000
Long-term capital losses (6,000 )

b. What is the Comers’ tax liability for 2020 if they report the following capital gains and losses for the year?

Short-term capital gains $ 1,500
Short-term capital losses 0
Long-term capital gains 13,000
Long-term capital losses (10,000 )

Schedule Y-1-Married Filing Jointly or Qualifying Widow(er)

If taxable income is over: But not over: The tax is:
$          0 $ 19,750 10% of taxable income
$ 19,750 $ 80,250 $1,975 plus 12% of the excess over $19,750
$ 80,250 $171,050 $9,235 plus 22% of the excess over $80,250
$171,050 $326,600 $29,211 plus 24% of the excess over $171,050
$326,600 $414,700 $66,543 plus 32% of the excess over $326,600
$414,700 $622,050 $94,735 plus 35% of the excess over $414,700
$622,050 $167,307.50 plus 37% of the excess over $622,050

Tax Rates for Net Capital Gains and Qualified Dividends

Rate* Taxable Income
Married Filing Jointly Married Filing Separately Single Head of Household Trusts and Estates
0% $0 - $80,000 $0 - $40,000 $0 - $40,000 $0 - $53,600 $0 - $2,650
15% $80,001 - $496,600 $40,001 - $248,300 $40,001 - $441,450 $53,601 - $469,050 $2,651 - $13,150
20% $496,601+ $248,301+ $441,451+ $469,051+ $13,151+

*This rate applies to the net capital gains and qualified dividends that fall within the range of taxable income specified in the table (net capital gains and qualified dividends are included in taxable income last for this purpose).
Answer #1

(a)

Particulars $
Salary ($68000+33000) 101000
Net STCG(STCG-STCL)=9000-2000 7000
Net LTCG(LTCG-LTCL)=15000-6000 9000
AGI 117,000
Less standard deduction 2020 (24,800)
Taxable income 92,200
Less prefrential income taxed separately LTCG (9000)
Income taxed at ordinary rates 83,200
Ordinary tax $9,235 plus 22% of the excess over $80,250=9235+2950*22% 9884
Preferential tax (9000*15% as taxable income is above the limit of $80000) 1350
Total tax liability(Ordinary+Preferential taxes) 11,234

(b)

Particulars $
Salary(Same as above) 101,000
Net STCG ($1500-$0) 1500
Net LTCG($13000-$10,000) 3000
AGI 105,500
Less Standard deduction (24,800)
Taxable income 80,700
Less preferential income taxed separately LTCG (3000)
Income taxable at ordinary rates 77,700
Ordinary tax $1,975 plus 12% of the excess over $19,750=1975+57950*12% 8929
Preferential Tax 3000*15% as taxable income exceeds specified limit of 80,000 which is 80700(net income including CG) 450
Total taxes 9379